# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ Appendix G Robustness to Out-of-Distribution data
In summary, the results in Tables 19, 21 and 23 show that our ICL approach is, in most cases, capable of robustly generalizing beyond its specific pre-training distribution when various aspects of this distribution are changed. While the performance sometimes decreases when a mismatch between training and testing data occurs, the drops in performance are almost always modest and, in many cases, almost negligible.
G.1 GLM scenarios
@@
Scenario 3 $\text{Laplace}(0,1)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 3.B $\text{Laplace}(0,\sqrt{2})$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 3.C $\text{Laplace}(0,2)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 5 $\text{Ga}(1,1)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 5.B $\text{Ga}(1,\sqrt{2})$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 5.C $\text{Ga}(1,2)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
@@ G.2 FA Scenarios
Scenario 1 $50$ $3$ $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,1)$ $3$
Scenario 1.B $50$ $3$ $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,2)$ $\mathcal{N}(0,1)$ $3$
Scenario 1.C $50$ $3$ $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,4)$ $\mathcal{N}(0,1)$ $3$
Scenario 2 $50$ $3$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,10)$ $\mathcal{N}(0,1)$ $3$
Scenario 2.B $50$ $3$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,10\cdot\sqrt{2})$ $\mathcal{N}(0,1)$ $3$
Scenario 2.C $50$ $3$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,20)$ $\mathcal{N}(0,1)$ $3$
Scenario 3 $25$ $5$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\mathcal{N}(0,1)$ $3$
Scenario 3 $25$ $5$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\cdot\sqrt{2})$ $\mathcal{N}(0,1)$ $3$
Scenario 3 $25$ $5$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,6)$ $\mathcal{N}(0,1)$ $3$
@@
Scenario 1 0.552 ($\pm$ 0.028) 0.034 (±($\pm$( ± 0.034) 0.289 ($\pm$ 0.083)
Scenario 1.B 0.826 ($\pm$ 0.066) 0.656 (±($\pm$( ± 0.384) 0.929 ($\pm$ 0.321)
Scenario 1.C 0.855 ($\pm$ 0.060) 0.837 (±($\pm$( ± 0.494) 1.135 ($\pm$ 0.461)
Scenario 2 0.542 ($\pm$ 0.006) 0.017 (±($\pm$( ± 0.006) 0.244 ($\pm$ 0.033)
Scenario 2.B 0.580 ($\pm$ 0.069) 0.087 (±($\pm$( ± 0.191) 0.393 ($\pm$ 0.291)
Scenario 2.C 0.589 ($\pm$ 0.076) 0.089 (±($\pm$( ± 0.113) 0.446 ($\pm$ 0.233)
Scenario 3 0.537 ($\pm$ 0.023) 0.024 (±($\pm$( ± 0.021) 0.259 ($\pm$ 0.088)
Scenario 3.B 0.544 ($\pm$ 0.028) 0.030 (±($\pm$( ± 0.021) 0.285 ($\pm$ 0.094)
Scenario 3.C 0.533 ($\pm$ 0.025) 0.021 (±($\pm$( ± 0.015) 0.347 ($\pm$ 0.152)
@@ G.3 GMM Scenarios
Scenario 2 $25$ $3$ $3$ $\text{Dir}(1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.B $25$ $3$ $3$ $\text{Dir}(0.5)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.C $25$ $3$ $3$ $\text{Dir}(2)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.D $25$ $3$ $3$ $\text{Dir}(1)$ $\text{IG}(5,2\cdot\sqrt{2})$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.E $25$ $3$ $3$ $\text{Dir}(1)$ $\text{IG}(5,4)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.F $25$ $3$ $3$ $\text{Dir}(1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,4\sigma^{2}_{i,m,l})$
Scenario 2.G $25$ $3$ $3$ $\text{Dir}(1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,5\sigma^{2}_{i,m,l})$
@@
Scenario 2 0.812 ($\pm$ 0.061) 0.159 (±($\pm$( ± 0.154) 2.314 ($\pm$ 0.926)
Scenario 2.B 0.829 ($\pm$ 0.050) 0.233 (±($\pm$( ± 0.161) 2.595 ($\pm$ 0.998)
Scenario 2.C 0.816 ($\pm$ 0.057) 0.149 (±($\pm$( ± 0.135) 2.272 ($\pm$ 0.654)
Scenario 2 0.812 ($\pm$ 0.061) 0.159 (±($\pm$( ± 0.154) 2.314 ($\pm$ 0.926)
Scenario 2.D 0.812 ($\pm$ 0.076) 0.148 (±($\pm$( ± 0.091) 2.557 ($\pm$ 0.837)
Scenario 2.E 0.880 ($\pm$ 0.057) 0.231 (±($\pm$( ± 0.109) 3.535 ($\pm$ 1.003)
Scenario 2 0.812 ($\pm$ 0.061) 0.159 (±($\pm$( ± 0.154) 2.314 ($\pm$ 0.926)
Scenario 2.F 0.821 ($\pm$ 0.076) 0.216 (±($\pm$( ± 0.214) 2.700 ($\pm$ 1.044
Scenario 2.G 0.844 ($\pm$ 0.046) 0.197 (±($\pm$( ± 0.124) 2.675 ($\pm$ 0.552)
@@
Appendix H Ablation: Using an MLP-based encoder
@@
Scenario 2 MLP 0.942 ($\pm$ 0.093) 1.783 (±($\pm$( ± 1.048) 2.503 ($\pm$ 0.814) 0.968 ($\pm$ 0.012) 1.528 (±($\pm$( ± 0.394) 2.271 ($\pm$ 0.315)
Transformer 0.839 ($\pm$ 0.072) 0.707 (±($\pm$( ± 0.658) 1.111 ($\pm$ 0.300) 0.768 ($\pm$ 0.033) 0.143 (±($\pm$( ± 0.089) 0.411 ($\pm$ 0.094)
Scenario 3 MLP 0.957 ($\pm$ 0.075) 2.236 (±($\pm$( ± 1.218) 2.681 ($\pm$ 1.130) 0.972 ($\pm$ 0.012) 1.658 (±($\pm$( ± 0.450) 2.076 ($\pm$ 0.427)
Transformer 0.611 ($\pm$ 0.070) 0.089 (±($\pm$( ± 0.114) 0.423 ($\pm$ 0.348) 0.576 ($\pm$ 0.027) 0.037 (±($\pm$( ± 0.026) 0.257 ($\pm$ 0.044)
Scenario 5 MLP 0.845 ($\pm$ 0.115) 1.066 (±($\pm$( ± 0.859) 1.166 ($\pm$ 0.996) 0.890 ($\pm$ 0.055) 1.223 (±($\pm$( ± 0.791) 1.102 ($\pm$ 0.383)
Transformer 0.621 ($\pm$ 0.063) 0.067 (±($\pm$( ± 0.080) 0.299 ($\pm$ 0.195) 0.610 ($\pm$ 0.045) 0.046 (±($\pm$( ± 0.020) 0.242 ($\pm$ 0.038)

### v2

@@ I.3 Gaussian Mixture Models
In summary, the results in Tables 31, 33 and 35 show that our ICL approach is, in most cases, capable of robustly generalizing beyond its specific pre-training distribution when various aspects of this distribution are changed. While the performance sometimes decreases when a mismatch between training and testing data occurs, the drops in performance are almost always modest and, in many cases, almost negligible.
N.1 GLM Scenarios
@@
Scenario 3 $\text{Laplace}(0,1)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 3.B $\text{Laplace}(0,\sqrt{2})$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 3.C $\text{Laplace}(0,2)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 5 $\text{Ga}(1,1)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 5.B $\text{Ga}(1,\sqrt{2})$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 5.C $\text{Ga}(1,2)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
@@
Scenario 1 $50$ 33 $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,1)$ 33
Scenario 1.B $50$ 33 $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,2)$ $\mathcal{N}(0,1)$ 33
Scenario 1.C $50$ 33 $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,4)$ $\mathcal{N}(0,1)$ 33
Scenario 2 $50$ 33 $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,10)$ $\mathcal{N}(0,1)$ 33
Scenario 2.B $50$ 33 $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,10\cdot\sqrt{2})$ $\mathcal{N}(0,1)$ 33
Scenario 2.C $50$ 33 $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,20)$ $\mathcal{N}(0,1)$ 33
Scenario 3 $25$ 55 $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\mathcal{N}(0,1)$ 33
Scenario 3 $25$ 55 $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\cdot\sqrt{2})$ $\mathcal{N}(0,1)$ 33
Scenario 3 $25$ 55 $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,6)$ $\mathcal{N}(0,1)$ 33
@@
Scenario 1 0.552 ($\pm$ 0.028) 0.034 $(\pm$ 0.034) 0.289 ($\pm$ 0.083)
Scenario 1.B 0.826 ($\pm$ 0.066) 0.656 $(\pm$ 0.384) 0.929 ($\pm$ 0.321)
Scenario 1.C 0.855 ($\pm$ 0.060) 0.837 $(\pm$ 0.494) 1.135 ($\pm$ 0.461)
Scenario 2 0.542 ($\pm$ 0.006) 0.017 $(\pm$ 0.006) 0.244 ($\pm$ 0.033)
Scenario 2.B 0.580 ($\pm$ 0.069) 0.087 $(\pm$ 0.191) 0.393 ($\pm$ 0.291)
Scenario 2.C 0.589 ($\pm$ 0.076) 0.089 $(\pm$ 0.113) 0.446 ($\pm$ 0.233)
Scenario 3 0.537 ($\pm$ 0.023) 0.024 $(\pm$ 0.021) 0.259 ($\pm$ 0.088)
Scenario 3.B 0.544 ($\pm$ 0.028) 0.030 $(\pm$ 0.021) 0.285 ($\pm$ 0.094)
Scenario 3.C 0.533 ($\pm$ 0.025) 0.021 $(\pm$ 0.015) 0.347 ($\pm$ 0.152)
@@
Scenario 2 $25$ 33 33 $\text{Dir}(1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.B $25$ 33 33 $\text{Dir}(0.5)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.C $25$ 33 33 $\text{Dir}(2)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.D $25$ 33 33 $\text{Dir}(1)$ $\text{IG}(5,2\cdot\sqrt{2})$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.E $25$ 33 33 $\text{Dir}(1)$ $\text{IG}(5,4)$ $\mathcal{N}(0,3\sigma^{2}_{i,m,l})$
Scenario 2.F $25$ 33 33 $\text{Dir}(1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,4\sigma^{2}_{i,m,l})$
Scenario 2.G $25$ 33 33 $\text{Dir}(1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,5\sigma^{2}_{i,m,l})$
@@
Scenario 2 0.812 ($\pm$ 0.061) 0.159 $(\pm$ 0.154) 2.314 ($\pm$ 0.926)
Scenario 2.B 0.829 ($\pm$ 0.050) 0.233 $(\pm$ 0.161) 2.595 ($\pm$ 0.998)
Scenario 2.C 0.816 ($\pm$ 0.057) 0.149 $(\pm$ 0.135) 2.272 ($\pm$ 0.654)
Scenario 2 0.812 ($\pm$ 0.061) 0.159 $(\pm$ 0.154) 2.314 ($\pm$ 0.926)
Scenario 2.D 0.812 ($\pm$ 0.076) 0.148 $(\pm$ 0.091) 2.557 ($\pm$ 0.837)
Scenario 2.E 0.880 ($\pm$ 0.057) 0.231 $(\pm$ 0.109) 3.535 ($\pm$ 1.003)
Scenario 2 0.812 ($\pm$ 0.061) 0.159 $(\pm$ 0.154) 2.314 ($\pm$ 0.926)
Scenario 2.F 0.821 ($\pm$ 0.076) 0.216 $(\pm$ 0.214) 2.700 ($\pm$ 1.044
Scenario 2.G 0.844 ($\pm$ 0.046) 0.197 $(\pm$ 0.124) 2.675 ($\pm$ 0.552)
@@
Appendix O Ablation: Dimensionality of the Problem
In this section, the effect of the dimensionality $K$ of the latent variable $\bm{z}\in\mathbb{R}^{K}$ for the GLM scenarios is investigated.
In summary, our results show that forward-KL based VI and in particular MAP solutions perform strongly in terms of predictive performance, which is in line with results first presented by Mittal et al. (2025b, a).
In table 36, we find that the advantages of the in-context learning approach to deteriorate for higher dimensionalities, with the variational inference methods using a Gaussian approximation performing well for 20 dimensions. This finding is line with work by Mittal et al. (2025b, a). For 50 dimensions we find that in many cases the used metrics do not allow to significantly discriminate the performance of the different approaches. Note that the randomness in the results, especially for higher dimensionalities, can in rare cases lead to better mean values. This is most likely not significant when taking the standard error into account.
Similar to scenarios 1,2 and 3, we find that in scenarios 4,5 and 6 (Table LABEL:tab:Dim_v2) the advantages of the in-context learning approach to deteriorate for higher dimensionalities, with the variational inference methods using a Gaussian approximation performing well for 20 dimensions. For 50 dimensions we find that in many cases the used metrics do not allow to significantly discriminate the performance of the different approaches. Note that the randomness in the results, especially for higher dimensionalities, can in rare cases lead to better mean values. This is most likely not significant when taking the standard error into account.
Finally, the results in Table 38 show that also for this scenario 7, the advantages of the in-context learning approach to deteriorate for higher dimensionalities. However, in this specific scenario the in-context learner is not significantly different from the VI methods in terms of C2ST and MMD for 20 dimensions. For 50 dimensions we find that the VI method using IAF performs well, together with the in-context learning approach in terms of MMD while the C2ST score does not indicate a clear winner and $\mathcal{W}_{2}$ favors the other methods. Note that the randomness in the results, especially for higher dimensionalities, can in rare cases lead to better mean values. This is most likely not significant when taking the standard error into account.
Table 36: Generalized Linear Models: Ablation with respect to the dimensionality of the problem on 50 synthetic and 17 real-world datasets for scenarios 1,2 and 3. All results within two standard errors of the best average result for each scenario are marked in bold. Due to the limitations of the number of features in the real-world data, we can only use 5 datasets for 20 and one dataset for 50 dimensions.
Scenario Dim. Model Synthetic Evaluation Real-World Evaluation
@@
Scenario 1    5 Laplace Approximation 1.000 ($\pm$ 0.000) 2.738 $(\pm$ 0.721) 0.825 ($\pm$ 0.279) 1.000 ($\pm$ 0.000) 2.150 $(\pm$ 0.323) 0.642 ($\pm$ 0.124)
VI: DiagonalNormal 0.904 ($\pm$ 0.076) 1.452 $(\pm$ 0.984) 0.669 ($\pm$ 0.301) 0.797 ($\pm$ 0.083) 0.612 $(\pm$ 0.511) 0.414 ($\pm$ 0.152)
VI: MultivariateNormal 0.750 ($\pm$ 0.128) 0.735 $(\pm$ 0.733) 0.565 ($\pm$ 0.292) 0.607 ($\pm$ 0.070) 0.167 $(\pm$ 0.196) 0.301 ($\pm$ 0.123)
VI: Structured Normal 0.753 ($\pm$ 0.126) 0.736 $(\pm$ 0.737) 0.570 ($\pm$ 0.310) 0.600 ($\pm$ 0.070) 0.169 $(\pm$ 0.214) 0.306 ($\pm$ 0.131)
VI: IAF 0.777 ($\pm$ 0.122) 0.864 $(\pm$ 0.844) 0.725 ($\pm$ 0.523) 0.683 ($\pm$ 0.132) 0.440 $(\pm$ 0.559) 0.503 ($\pm$ 0.383)
HMC 0.745 ($\pm$ 0.130) 0.722 $(\pm$ 0.732) 0.569 ($\pm$ 0.301) 0.595 ($\pm$ 0.075) 0.173 $(\pm$ 0.213) 0.321 ($\pm$ 0.140)
ICL (ours) 0.765 ($\pm$ 0.123) 0.767 $(\pm$ 0.727) 0.585 ($\pm$ 0.301) 0.614 ($\pm$ 0.074) 0.175 $(\pm$ 0.219) 0.310 ($\pm$ 0.138)
Scenario 1    20 Laplace Approximation 1.000 ($\pm$ 0.000) 2.237 ($\pm$ 0.024) 3.252 ($\pm$ 1.172) 1.000 ($\pm$ 0.000) 2.206 ($\pm$ 0.021) 2.792 ($\pm$ 0.339)
VI: DiagonalNormal 0.843 ($\pm$ 0.204) 1.056 ($\pm$ 0.869) 2.976 ($\pm$ 0.927) 0.983 ($\pm$ 0.019) 1.217 ($\pm$ 0.463) 2.406 ($\pm$ 0.348)
VI: MultivariateNormal 0.789 ($\pm$ 0.140) 0.714 ($\pm$ 0.351) 2.774 ($\pm$ 0.995) 0.768 ($\pm$ 0.120) 0.180 ($\pm$ 0.159) 2.064 ($\pm$ 0.306)
VI: Structured Normal 0.792 ($\pm$ 0.109) 0.708 ($\pm$ 0.153) 2.703 ($\pm$ 1.069) 0.668 ($\pm$ 0.080) 0.168 ($\pm$ 0.071) 2.052 ($\pm$ 0.275)
VI: IAF 0.832 ($\pm$ 0.196) 0.847 ($\pm$ 1.016) 4.015 ($\pm$ 0.415) 0.939 ($\pm$ 0.024) 0.508 ($\pm$ 0.200) 3.140 ($\pm$ 0.290)
ICL (ours) 0.849 ($\pm$ 0.171) 0.844 ($\pm$ 0.905) 4.564 ($\pm$ 0.622) 0.970 ($\pm$ 0.020) 0.724 ($\pm$ 0.287) 4.250 ($\pm$ 0.312)
Scenario 1    50 Laplace Approximation 1.000 ($\pm$ 0.000) 2.401 ($\pm$ 0.282) 5.152 ($\pm$ 2.268) 1.000 ($\pm$ nan) 2.339 ($\pm$ nan) 6.642 ($\pm$ nan)
VI: DiagonalNormal 0.812 ($\pm$ 0.197) 0.956 ($\pm$ 1.016) 5.541 ($\pm$ 2.356) 0.915 ($\pm$ nan) 0.508 ($\pm$ nan) 6.200 ($\pm$ nan)
VI: MultivariateNormal 0.839 ($\pm$ 0.148) 0.926 ($\pm$ 0.682) 5.514 ($\pm$ 2.370) 0.905 ($\pm$ nan) 0.790 ($\pm$ nan) 6.258 ($\pm$ nan)
VI: Structured Normal 0.823 ($\pm$ 0.160) 0.844 ($\pm$ 0.480) 5.752 ($\pm$ 2.098) 0.910 ($\pm$ nan) 1.122 ($\pm$ nan) 6.898 ($\pm$ nan)
VI: IAF 0.820 ($\pm$ 0.182) 0.814 ($\pm$ 0.987) 6.696 ($\pm$ 1.207) 0.938 ($\pm$ nan) 0.256 ($\pm$ nan) 6.869 ($\pm$ nan)
ICL (ours) 0.787 ($\pm$ 0.217) 1.015 ($\pm$ 1.255) 8.278 ($\pm$ 0.821) 0.979 ($\pm$ nan) 0.413 ($\pm$ nan) 8.368 ($\pm$ nan)
Scenario 2    5 Laplace Approximation 1.000 ($\pm$ 0.000) 4.853 $(\pm$ 2.333) 5.770 ($\pm$ 5.946) 1.000 ($\pm$ 0.000) 2.572 $(\pm$ 0.206) 0.809 ($\pm$ 0.149)
VI: DiagonalNormal 0.957 ($\pm$ 0.091) 3.906 $(\pm$ 2.679) 5.628 ($\pm$ 6.092) 0.892 ($\pm$ 0.044) 0.847 $(\pm$ 0.389) 0.530 ($\pm$ 0.175)
VI: MultivariateNormal 0.910 ($\pm$ 0.131) 3.407 $(\pm$ 2.781) 5.584 ($\pm$ 6.104) 0.820 ($\pm$ 0.031) 0.243 $(\pm$ 0.148) 0.408 ($\pm$ 0.118)
VI: Structured Normal 0.908 ($\pm$ 0.119) 3.139 $(\pm$ 2.763) 5.480 ($\pm$ 6.164) 0.824 ($\pm$ 0.023) 0.215 $(\pm$ 0.110) 0.392 ($\pm$ 0.109)
VI: IAF 0.968 ($\pm$ 0.063) 4.416 $(\pm$ 2.473) 7.474 ($\pm$ 6.235) 0.888 ($\pm$ 0.067) 0.921 $(\pm$ 0.860) 0.942 ($\pm$ 0.733)
ICL (ours) 0.839 ($\pm$ 0.072) 0.707 $(\pm$ 0.658) 1.111 ($\pm$ 0.300) 0.768 ($\pm$ 0.033) 0.143 $(\pm$ 0.089) 0.411 ($\pm$ 0.094)
Scenario 2    20 Laplace Approximation 1.000 ($\pm$ 0.000) 2.314 ($\pm$ 0.237) 3.069 ($\pm$ 1.168) 1.000 ($\pm$ 0.000) 2.222 ($\pm$ 0.018) 2.847 ($\pm$ 0.305)
VI: DiagonalNormal 0.904 ($\pm$ 0.168) 1.292 ($\pm$ 0.937) 2.863 ($\pm$ 0.919) 0.990 ($\pm$ 0.009) 1.277 ($\pm$ 0.452) 2.483 ($\pm$ 0.318)
VI: MultivariateNormal 0.851 ($\pm$ 0.134) 0.492 ($\pm$ 0.547) 2.694 ($\pm$ 0.916) 0.843 ($\pm$ 0.069) 0.243 ($\pm$ 0.170) 2.166 ($\pm$ 0.266)
VI: Structured Normal 0.697 ($\pm$ 0.065) 0.070 ($\pm$ 0.099) 2.497 ($\pm$ 0.993) 0.655 ($\pm$ 0.031) 0.029 ($\pm$ 0.025) 2.191 ($\pm$ 0.271)
VI: IAF 0.916 ($\pm$ 0.110) 1.062 ($\pm$ 1.076) 4.191 ($\pm$ 0.623) 0.952 ($\pm$ 0.025) 0.515 ($\pm$ 0.242) 3.331 ($\pm$ 0.371)
ICL (ours) 0.955 ($\pm$ 0.057) 1.131 ($\pm$ 1.035) 4.945 ($\pm$ 0.836) 0.968 ($\pm$ 0.020) 0.724 ($\pm$ 0.278) 4.356 ($\pm$ 0.302)
Scenario 2    50 Laplace Approximation 1.000 ($\pm$ 0.000) 2.437 ($\pm$ 0.271) 5.728 ($\pm$ 1.358) 1.000 ($\pm$ nan) 2.350 ($\pm$ nan) 5.620 ($\pm$ nan)
VI: DiagonalNormal 0.853 ($\pm$ 0.182) 0.787 ($\pm$ 0.687) 6.224 ($\pm$ 1.225) 0.996 ($\pm$ nan) 1.080 ($\pm$ nan) 5.426 ($\pm$ nan)
VI: MultivariateNormal 0.878 ($\pm$ 0.150) 0.688 ($\pm$ 0.620) 6.206 ($\pm$ 1.244) 0.994 ($\pm$ nan) 0.791 ($\pm$ nan) 5.305 ($\pm$ nan)
VI: Structured Normal 0.865 ($\pm$ 0.081) 0.186 ($\pm$ 0.169) 5.874 ($\pm$ 1.233) 0.819 ($\pm$ nan) 0.093 ($\pm$ nan) 5.660 ($\pm$ nan)
VI: IAF 0.909 ($\pm$ 0.130) 0.649 ($\pm$ 0.650) 7.465 ($\pm$ 0.335) 0.985 ($\pm$ nan) 0.426 ($\pm$ nan) 6.426 ($\pm$ nan)
ICL (ours) 0.972 ($\pm$ 0.039) 0.741 ($\pm$ 0.713) 8.313 ($\pm$ 0.608) 0.971 ($\pm$ nan) 0.405 ($\pm$ nan) 7.718 ($\pm$ nan)
Scenario 3    5 Laplace Approximation 1.000 ($\pm$ 0.000) 2.203 ($\pm$ 0.997) 1.170 ($\pm$ 0.949) 1.000 ($\pm$ 0.000) 1.841 ($\pm$ 0.185) 0.729 ($\pm$ 0.175)
VI: DiagonalNormal 0.866 ($\pm$ 0.101) 1.069 ($\pm$ 1.150) 0.846 ($\pm$ 0.747) 0.797 ($\pm$ 0.083) 0.526 ($\pm$ 0.361) 0.480 ($\pm$ 0.207)
VI: MultivariateNormal 0.656 ($\pm$ 0.131) 0.445 ($\pm$ 1.061) 0.660 ($\pm$ 0.737) 0.560 ($\pm$ 0.035) 0.032 ($\pm$ 0.028) 0.249 ($\pm$ 0.069)
VI: Structured Normal 0.653 ($\pm$ 0.125) 0.421 ($\pm$ 0.993) 0.659 ($\pm$ 0.736) 0.552 ($\pm$ 0.028) 0.027 ($\pm$ 0.015) 0.239 ($\pm$ 0.055)
VI: IAF 0.751 ($\pm$ 0.148) 0.939 ($\pm$ 1.349) 0.964 ($\pm$ 0.924) 0.673 ($\pm$ 0.141) 0.399 ($\pm$ 0.543) 0.563 ($\pm$ 0.433)
ICL (ours) 0.611 ($\pm$ 0.070) 0.089 ($\pm$ 0.114) 0.423 ($\pm$ 0.348) 0.576 ($\pm$ 0.027) 0.037 ($\pm$ 0.026) 0.257 ($\pm$ 0.044)
Scenario 3    20 Laplace Approximation 1.000 ($\pm$ 0.000) 2.726 ($\pm$ 1.116) 4.127 ($\pm$ 1.927) 1.000 ($\pm$ 0.000) 2.234 ($\pm$ 0.092) 3.589 ($\pm$ 0.519)
VI: DiagonalNormal 0.912 ($\pm$ 0.134) 1.704 ($\pm$ 1.467) 3.933 ($\pm$ 1.574) 0.983 ($\pm$ 0.014) 1.298 ($\pm$ 0.443) 3.147 ($\pm$ 0.557)
VI: MultivariateNormal 0.863 ($\pm$ 0.113) 0.937 ($\pm$ 1.174) 3.754 ($\pm$ 1.650) 0.796 ($\pm$ 0.099) 0.268 ($\pm$ 0.226) 2.645 ($\pm$ 0.466)
VI: Structured Normal 0.768 ($\pm$ 0.109) 0.302 ($\pm$ 0.518) 3.151 ($\pm$ 1.663) 0.722 ($\pm$ 0.073) 0.131 ($\pm$ 0.141) 2.579 ($\pm$ 0.399)
VI: IAF 0.908 ($\pm$ 0.133) 1.657 ($\pm$ 1.476) 5.543 ($\pm$ 1.120) 0.936 ($\pm$ 0.041) 0.548 ($\pm$ 0.341) 3.678 ($\pm$ 0.670)
ICL (ours) 0.902 ($\pm$ 0.076) 1.053 ($\pm$ 0.782) 6.206 ($\pm$ 0.783) 0.932 ($\pm$ 0.019) 0.635 ($\pm$ 0.183) 5.281 ($\pm$ 0.317)
Scenario 3    50 Laplace Approximation 1.000 ($\pm$ 0.000) 2.700 ($\pm$ 0.789) 8.841 ($\pm$ 1.691) 1.000 ($\pm$ nan) 2.348 ($\pm$ nan) 7.049 ($\pm$ nan)
VI: DiagonalNormal 0.870 ($\pm$ 0.127) 1.154 ($\pm$ 1.321) 9.180 ($\pm$ 1.513) 0.997 ($\pm$ nan) 1.393 ($\pm$ nan) 6.791 ($\pm$ nan)
VI: MultivariateNormal 0.896 ($\pm$ 0.101) 1.027 ($\pm$ 1.157) 9.175 ($\pm$ 1.555) 0.998 ($\pm$ nan) 1.092 ($\pm$ nan) 6.667 ($\pm$ nan)
VI: Structured Normal 0.873 ($\pm$ 0.112) 0.539 ($\pm$ 0.667) 9.118 ($\pm$ 1.538) 0.958 ($\pm$ nan) 0.420 ($\pm$ nan) 6.665 ($\pm$ nan)
VI: IAF 0.869 ($\pm$ 0.124) 0.751 ($\pm$ 0.939) 9.917 ($\pm$ 0.870) 0.971 ($\pm$ nan) 0.417 ($\pm$ nan) 7.411 ($\pm$ nan)
ICL (ours) 0.931 ($\pm$ 0.062) 0.784 ($\pm$ 0.884) 10.063 ($\pm$ 0.930) 0.965 ($\pm$ nan) 0.347 ($\pm$ nan) 8.482 ($\pm$ nan)
Table 37: Generalized Linear Models: Ablation with respect to the dimensionality of the problem on 50 synthetic and 17 real-world datasets for scenarios 4, 5 and 6. All results within two standard errors of the best average result for each scenario are marked in bold. Due to the limitations of the number of features in the real-world data, we can only use 5 datasets for 20 and one dataset for 50 dimensions.
Scenario Dim. Model Synthetic Evaluation Real-World Evaluation